hello i submitted these questions and get answer but what i needed to know was HOW do we get the answers I want to see the work so i learn, so i am resubmitting hoping for this:
I have been working on this all week and I just cannot get it. I know this is short notice but it is due with in the next 12 hours PLEASE HELP!
A plasma TV company determines that its total profit on the production and sale of x TVs is given by: P(x)= -x^2+812x-9600
the question that i am stuck on are as follows, and i have to show the work:
1. Find the solutions and explain what these solutions represent.
2. How many TV�s must be produced to reach a minimum or maximum profit?
3.What is this minimum or maximum profit?
4.What is the vertex of this graph? How is this point related to your answers from e and f?
I also have to graph this function. Include solutions (roots), vertex, axis of symmetry, and labels for each.

Thanks in advance

Answer by scott8148(6628)

1. find the roots by setting P(x) equal to zero ___ -x^2 + 812x - 9600 = 0
___ dividing by -1 ___ x^2 - 812x + 9600 = 0
___ factoring ___ (x - 12)(x - 800) = 0
___ the roots are 12 and 800
___ the lower number represents the break even point where the costs equal the sales
___ the higher number probably represents market saturation (I do math, not finance)
to much product for the market and the price declines (and/or sales drop), causing profits to fall

2. the min (or max, in this case) lies midway between the roots (their average)
___ (800 + 12) / 2 = 406

3. plug 406 into the profit function ___ P(406) = -406^2 + 812(406) - 9600

4. the x-coordinate of the vertex is 406 , the y-coordinate is the answer from #3
___ this represents the max profit and the number of sets needed to achieve it


the axis of symmetry passes through the vertex
___ it is a vertical line whose equation is ___ x = 406


the parabola opens downward
it passes through the roots and the vertex (which is the "nose" of the parabola)
the axis of symmetry is the midline of the parabola